# Imaging Signatures of the Israel Junction: Photon Ring Evolution in Dynamical Thin Shell Schwarzschild Spacetimes

### v1

@@ I Introduction
In general relativity, joining two spacetimes along a hypersurface is a foundational and important problem. Many physical situations naturally exhibit geometry that is defined in pieces: models of stellar collapse often describe a fluid interior joined to an exterior Schwarzschild region; accretion flows and cosmological phase transitions are likewise idealized as two spacetime regions matched at a common hypersurface, and so on. Since Darmois’s early systematic treatment of non-singular matchings, it has been clarified that, on a timelike or spacelike matching hypersurface $\Sigma$, the absence of $\delta$-type singularities in the curvature requires continuity of both the induced metric and the extrinsic curvature, namely the Darmois–Lichnerowicz conditions MSM_1927__25__1_0; 1955trdl.book.....L. Subsequently, within a distributional framework that allows a $\delta$-source on $\Sigma$, Israel established the correspondence between the jump of the extrinsic curvature and the surface stress–energy tensor, leading to the now standard Israel junction conditions Israel:1966rt:
@@
which interpret geometric discontinuities as a thin shell source localized on the hypersurface; extensions to null hypersurfaces have also been developed Barrabes:1991ng.
@@
Since their introduction, the Israel junction conditions have seen broad applications in gravitational physics. In the collapse model by Oppenheimer and Snyder, one glues a pressureless dust interior to an exterior Schwarzschild spacetime along the stellar surface in order to track the boundary evolution and demonstrate horizon formation Oppenheimer:1939ue. In early thin shell wormhole constructions, one “cuts and pastes” two Schwarzschild exteriors, localizing matter on the shell and using the junction conditions to relate the surface stress-energy to the throat geometry Visser:1989kg; Poisson:1995sv. In the gravastar scenario, a Schwarzschild exterior is joined to a de Sitter interior by a thin shell. From the junction conditions, an effective potential for the shell radius can be derived. The analysis of this potential shows that stable, horizonless compact configurations can exist for suitable equations of state and parameter choices Mazur:2004fk; Mazur:2001fv; Visser:2003ge. Finally, one may glue a flat interior to an exterior Reissner-Nordström spacetime by means of a charged thin shell to probe weak cosmic censorship Boulware:1973tlq.
@@
On the observational side, black hole physics is transitioning from pure theoretical study to quantitative tests. Since the Event Horizon Telescope (EHT) released the first image of M87* in 2019 and of Sgr A* in 2022, tests at event horizon scales have become possible EventHorizonTelescope:2019dse; EventHorizonTelescope:2022wkp. Within general relativity, the Schwarzschild solution exhibits a distinctive geometric feature: an unstable spherical photon orbit outside the horizon, the photon sphere, which sets the critical boundary of the observed shadow and governs the strong-deflection lensing features and the sequence of higher order photon rings produced by multiple windings of light Virbhadra:1999nm; Johnson:2019ljv. Thus, studies of photon spheres are key probes of general relativity in the strong field regime Lupsasca:2024xhq.
@@
Research on black hole images has a long history. Early studies largely concentrated on characterizing the critical curves of the Schwarzschild and Kerr spacetimes Synge:1966okc; Bardeen:1973tla, that is, the images governed by the photon sphere. These curves demarcate bright and dark regions on the observer’s screen, and this program has been gradually extended to a variety of gravitational theories. With improving observational capabilities, theoretical models have begun to systematically include emission from accretion flows, redshift effects, polarization of light, and related processes Gralla:2019xty; Hou:2024qqo. As a result, the black hole image is no longer determined solely by the photon sphere, but by the combined effects of spacetime lensing and various astrophysical processes Hou:2022eev; Meng:2025ivb; Cao:2023par; Cao:2024kht; Li:2025wmd. Even so, many existing works still adopt stationary black holes. In reality, black holes evolve due to collapse, perturbations, or sustained accretion. When imaging such dynamical spacetimes, one must take into account light travel time delays caused by differences in path length. This requirement makes theoretical models closer to actual observations and is becoming increasingly important. In recent years, studies of imaging in dynamical black hole spacetimes have continued to emerge zhang2025emergingblackholeshadow; Liang_2025; wang2025dynamicshadowblackhole.
@@
Motivated by the above considerations, we examine the observable appearance of black holes endowed with a thin spherical shell and aim to identify features that may arise in spacetimes constructed by junctions. Thin shell spacetimes are concrete realizations of the Israel junction conditions, and by studying their imaging characteristics we can indirectly probe whether Israel junction geometries may be realized in strong gravity environments and place observational constraints on such models. To the best of our knowledge, most previous imaging studies of spacetimes built from junctions have focused on wormholes Nedkova:2013msa; Ohgami:2015nra; Peng:2021osd; Macedo:2025ipc. Although black holes with shells have been considered, existing works typically model the spherical shell as accreting matter and assess how it modifies otherwise standard static black hole images Narayan:2019imo; Wu:2024juj. In contrast, we investigate how matching distinct Schwarzschild geometries across a spherical shell modifies the spacetime geometry itself, and we analyze the resulting images for both static and dynamical shells.
@@ II Thin shell junction of two Schwarzschild spacetimes
In this section, we match two Schwarzschild spacetimes by a spherical shell, derive the equation of motion for the shell, and analyze its dynamics. There are already many works on thin shell junctions PhysRevD.44.1891; Gao:2008jy and on light propagation in such joined spacetimes Wang:2020emr. In this and the next section, we mainly give a relatively complete but still brief review of these results in preparation for the later image analysis.
@@
By the Israel junction conditions Israel:1966rt, the induced metric on the matching hypersurface $\Sigma$ must be continuous, $h_{ab}]=0$, where $X]=X_{+}-X_{-}$, whereas its normal derivative may be discontinuous. The resulting jump in the extrinsic curvature, $K_{ab}]$, encodes a surface stress-energy tensor localized on $\Sigma$.
@@
So far we have only required continuity of the induced metric along the shell. The expressions in Eq. (7) are computed separately in the two bulks and, a priori, need not represent the same geometric vector on $\Sigma$. However, if the unit normals from the two sides are not identified as a single normal field on $\Sigma$, the jump $K_{ab}]$ and the associated distributional definitions of the connection and curvature become ambiguous. This issue is more pronounced when three or more bulk spacetimes are glued along a single hypersurface. See Ref. Shen:2024dun and Appendix A for a detailed discussion. To ensure a well-defined differentiable structure and distributional curvature at the shell, we therefore identify the unit normals on the two sides at $\Sigma$:
@@
We consider a pressureless thin shell with surface stress-energy tensor $S^{i}{}_{j}=\mathrm{diag}(-\sigma,0,0)$. From the Israel conditions Israel:1966rt
@@
Eq. (16) shows that $m_{\mathrm{sh}}$ is constant. Using Eq. (15) with $m_{\mathrm{sh}}=\mathrm{const}$ yields Wang:2020emr
@@
Figure 1: Schematic effective potential $V(R)$ for the thin spherical shell. The solid black curve shows $V(R)$; the horizontal dashed lines in red, blue, and green mark $e^{2}$ corresponding to the cases $e^{2}\!<\!1$, $e^{2}\!=\!1$, and $e^{2}\!>\!1$, respectively. For $e^{2}<1$, the shell undergoes bound motion: it can reach at most the turning point $R_{\max}$, i.e., the intersection of the red dashed line with the black curve $V(R)$.
@@
Requiring $\sigma\geq 0$ implies $e\geq 0$. For $R>0$ the effective potential is monotonically increasing and concave, i.e. $V^{\prime}(R)>0$ and $V^{\prime\prime}(R)<0$. Fig. 1 schematically illustrates $V(R)$. When $e>1$ the motion is unbound and the shell can expand to spatial infinity; when $e=1$ the motion is critical and the shell reaches infinity with vanishing asymptotic velocity; when $e<1$ there is a turning point $R_{\max}$ defined by $V(R_{\max})=e^{2}$. Consequently, a pressureless shell admits no static, stable configuration, and static stability therefore requires nonzero pressure. Furthermore, any static shell must satisfy the necessary geometric constraint $R\geq 3m_{-}$, i.e., it must lie outside the interior photon sphere PhysRevD.44.1891.
@@ III Null geodesics in the joined spacetime
$\tan\alpha=\frac{r\,k^{\phi}}{|k^{r}|}=\frac{L}{r\,|k^{r}|},$ (26)
@@
$\tan\alpha_{+}=\frac{f_{-}(R)\,\tan\alpha_{-}}{\ \beta_{+}(R)\beta_{-}(R)-\dot{R}^{\,2}-s\dot{R}\bigl(\beta_{+}(R)-\beta_{-}(R)\bigr)\,\sqrt{\,1+f_{-}(R)\tan^{2}\!\alpha_{-}\,}\ }.$ (27)
@@
$E_{+}=\sqrt{\frac{f_{+}(R)}{f_{-}(R)}}\,E_{-},\qquad b_{+}=\sqrt{\frac{f_{-}(R)}{f_{+}(R)}}\,b_{-},\qquad\tan\alpha_{+}=\sqrt{\frac{f_{-}(R)}{f_{+}(R)}}\,\tan\alpha_{-}\;.$ (28)
Thus, for a static shell, a photon exiting the shell is redshifted and its refraction angle increases.
We now clarify what we mean by the “refraction angle”. We do not refer to the incident angle measured in an orthonormal frame comoving with the shell. Since the photon four momentum is continuous as a geometric vector, the direction measured in that local frame does not jump across the shell. Instead, our “refraction angle” is defined with respect to the static observers in the inner and outer Schwarzschild regions. Owing to the difference in the Killing fields across the shell, the static observers on either side are defined differently. As Eq. (11) shows, their orthonormal bases do not match at the shell. Consequently, observer-dependent quantities such as the refraction angle and frequency undergo a change across the shell. From Eq. (28) one sees that, in the static case, $\sqrt{f}\,\tan\alpha$ is invariant across the shell, which suggests interpreting $\sqrt{f}$ as an effective “gravitational refractive index”.
The changes in photon energy when light crosses the shell are encoded in the redshift factor, which directly influences the final image. For collapsing shells, see Ref. cbzn-pzrw for a systematic treatment; here we briefly summarize and adapt the relevant results. Define $g=\omega_{\text{obs}}/\omega_{\text{em}}$. In a static, spherically symmetric background, frequencies are defined with respect to static observers, then
@@ IV Static Shell Imaging
Having completed the junction of two Schwarzschild spacetimes and the analysis of null geodesics, we now turn to the imaging features of the resulting joined geometry. We begin with a static spherical shell for two reasons. First, the static case is simpler and allows us to identify the concrete impact of the shell on the image and the underlying mechanism, thereby laying the groundwork for understanding the various effects present for a dynamical shell. Second, the geometry obtained by matching two spacetimes along a static thin shell is of practical interest in its own right. For example, static shells are sometimes employed as idealized models of dark matter distributions around black holes Adhikari:2014lna; Qian:2021qow.
@@
where $r_{h_{-}}$ is the event horizon radius of the inner Schwarzschild region. We adopt this model precisely because our interest lies in image features controlled by the photon sphere. For a geometrically thin and optically thin disk, the photon ring of the final image is largely insensitive to the detailed radial emissivity, and extending the disk to the inner Schwarzschild horizon ensures that all trajectories relevant to the ring are included. For further discussion and comparisons of accretion disk models, see our previous work Cao:2023par; Cao:2024kht; Li:2025wmd. Accounting for the redshift factor, the observed specific intensity reads
@@
The V-shaped structure in Fig. 3(b) directly reflects the shell’s refraction effect. As derived in Sec. III, rays refract when traversing the shell, altering their propagation direction. The left panel of Fig. 4 shows rays that intersect the disk twice. The shell refracts rays with small $b$ outward, which increases the radius $r$ of the second intersection with the disk. As a result, $r(b)$ becomes nonmonotonic near $b\approx 3\sqrt{3}$ and takes on a V-shape in the plot. Similar V-shaped features of the transfer function $r(b)$ have, to our knowledge, been reported mainly in spacetimes with naked singularities or wormholes Luo:2023wru; Rueda:2025rlj; Tan:2025cte. In contrast, in standard black hole spacetimes without a refractive shell, $r(b)$ is typically monotonic. In our setup, we find that for suitable shell parameters, adding a refractive shell around a black hole can produce such a V-shaped structure in the transfer function. The same behavior occurs for rays that intersect the disk three times. This is the reason for the $\Lambda$-shaped peak in Fig. 3(a).
(a) $m_{-}=0.1,R_{sh}=5$
(b) $m_{-}=0.1,R_{sh}=3.1$
(c) $m_{-}=0.1,R_{sh}=2.9$
(d) $m_{-}=0.1,R_{sh}=2.1$
Figure 5: intensity profiles and transfer function $r(b)$ for several shell radii $R_{\rm sh}$ at fixed $m_{-}=0.1$.
(a) $m_{-}=0.1,R_{sh}=5$
(b) $m_{-}=0.1,R_{sh}=3.1$
(c) $m_{-}=0.1,R_{sh}=2.9$
(d) $m_{-}=0.1,R_{sh}=2.1$
Figure 6: Black hole images for various shell radii $R_{\rm sh}$ at $m_{-}=0.1$. In panel (d), we use a different intensity scale in order to highlight the double photon ring structure.
@@
(a) $m_{-}=0.8,R_{sh}=5$
(b) $m_{-}=0.8,R_{sh}=3.1$
(c) $m_{-}=0.8,R_{sh}=2.9$
(d) $m_{-}=0.8,R_{sh}=2.6$
Figure 7: intensity profiles and transfer function $r(b)$ for several shell radii $R_{\rm sh}$ at fixed $m_{-}=0.8$.
(a) $m_{-}=0.8,R_{sh}=5$
(b) $m_{-}=0.8,R_{sh}=3.1$
(c) $m_{-}=0.8,R_{sh}=2.9$
(d) $m_{-}=0.8,R_{sh}=2.6$
Figure 8: Black hole images for various shell radii $R_{\rm sh}$ at $m_{-}=0.8$.

### v2

@@ I Introduction
In general relativity, joining two spacetimes along a hypersurface is a foundational and important problem. Many physical situations naturally exhibit geometry that is defined in pieces: models of stellar collapse often describe a fluid interior joined to an exterior Schwarzschild region; accretion flows and cosmological phase transitions are likewise idealized as two spacetime regions matched at a common hypersurface, and so on. Since Darmois’s early systematic treatment of non-singular matchings, it has been clarified that, on a timelike or spacelike matching hypersurface $\Sigma$, the absence of $\delta$-type singularities in the curvature requires continuity of both the induced metric and the extrinsic curvature, namely the Darmois–Lichnerowicz conditions Darmois (1927); Lichnerowicz (1955). Subsequently, within a distributional framework that allows a $\delta$-source on $\Sigma$, Israel established the correspondence between the jump of the extrinsic curvature and the surface stress–energy tensor, leading to the now standard Israel junction conditions Israel (1966):
@@
which interpret geometric discontinuities as a thin shell source localized on the hypersurface; extensions to null hypersurfaces have also been developed Barrabes and Israel (1991).
@@
Since their introduction, the Israel junction conditions have seen broad applications in gravitational physics. In the collapse model by Oppenheimer and Snyder, one glues a pressureless dust interior to an exterior Schwarzschild spacetime along the stellar surface in order to track the boundary evolution and demonstrate horizon formation Oppenheimer and Snyder (1939). In early thin shell wormhole constructions, one “cuts and pastes” two Schwarzschild exteriors, localizing matter on the shell and using the junction conditions to relate the surface stress-energy to the throat geometry Visser (1989); Poisson and Visser (1995). In the gravastar scenario, a Schwarzschild exterior is joined to a de Sitter interior by a thin shell. From the junction conditions, an effective potential for the shell radius can be derived. The analysis of this potential shows that stable, horizonless compact configurations can exist for suitable equations of state and parameter choices Mazur and Mottola (2004, 2023); Visser and Wiltshire (2004). Finally, one may glue a flat interior to an exterior Reissner-Nordström spacetime by means of a charged thin shell to probe weak cosmic censorship Boulware (1973). Beyond these representative examples, the Israel junction conditions have found widespread application across a broad range of contexts in gravitational theory Tsukamoto (2021); Capozziello et al. (2024); De Bianchi et al. (2025); Capozziello et al. (2025). More generally, thin shell models can serve as useful idealizations in a variety of astrophysical settings. For a static shell, they can approximate a thin layer of dark matter around a black hole Adhikari et al. (2014); Qian et al. (2022). For a dynamical collapsing shell, they can describe late-stage fallback of matter during collapse and spherically symmetric accretion Liu and Zhang (2009).
@@
On the observational side, black hole physics is transitioning from pure theoretical study to quantitative tests. Since the Event Horizon Telescope (EHT) released the first image of M87* in 2019 and of Sgr A* in 2022, tests at event horizon scales have become possible Akiyama and others (2019, 2022); Mizuno et al. (2018). Within general relativity, the Schwarzschild solution exhibits a distinctive geometric feature: an unstable spherical photon orbit outside the horizon, the photon sphere, which sets the critical boundary of the observed shadow and governs the strong-deflection lensing features and the sequence of higher order photon rings produced by multiple windings of light Virbhadra and Ellis (2000); Johnson and others (2020); Vagnozzi and others (2023). Thus, studies of photon spheres are key probes of general relativity in the strong field regime Lupsasca et al. (2024); Di Filippo and Rezzolla (2025).
@@
Research on black hole images has a long history. Early studies largely concentrated on characterizing the critical curves of the Schwarzschild and Kerr spacetimes Synge (1966); Bardeen (1973), that is, the images governed by the photon sphere. These curves demarcate bright and dark regions on the observer’s screen, and this program has been gradually extended to a variety of gravitational theories. With improving observational capabilities, theoretical models have begun to systematically include emission from accretion flows, redshift effects, polarization of light, and related processes Gralla et al. (2019); Hou et al. (2025). As a result, the black hole image is no longer determined solely by the photon sphere, but by the combined effects of spacetime lensing and various astrophysical processes Hou et al. (2022); Meng et al. (2025); Cao et al. (2024b, a); Li et al. (2025); Uniyal et al. (2025). Even so, many existing works still adopt stationary black holes. In reality, black holes evolve due to collapse, perturbations, or sustained accretion. When imaging such dynamical spacetimes, one must take into account light travel time delays caused by differences in path length. This requirement makes theoretical models closer to actual observations and is becoming increasingly important. In recent years, studies of imaging in dynamical black hole spacetimes have continued to emerge Zhang et al. (2025); Liang et al. (2025); Wang et al. (2025).
@@
Motivated by the above considerations, we examine the observable appearance of black holes endowed with a thin spherical shell and aim to identify features that may arise in spacetimes constructed by junctions. Thin shell spacetimes are concrete realizations of the Israel junction conditions, and by studying their imaging characteristics we can indirectly probe whether Israel junction geometries may be realized in strong gravity environments and place observational constraints on such models. To the best of our knowledge, most previous imaging studies of spacetimes built from junctions have focused on wormholes Nedkova et al. (2013); Ohgami and Sakai (2015); Peng et al. (2021); Macedo et al. (2025). Although black holes with shells have been considered, existing works typically model the spherical shell as accreting matter and assess how it modifies otherwise standard static black hole images Narayan et al. (2019); Wu et al. (2025). In contrast, we investigate how matching distinct Schwarzschild geometries across a spherical shell modifies the spacetime geometry itself, and we analyze the resulting images for both static and dynamical shells.
@@ II Thin shell junction of two Schwarzschild spacetimes
In this section, we match two Schwarzschild spacetimes by a spherical shell, derive the equation of motion for the shell, and analyze its dynamics. There are already many works on thin shell junctions Brady et al. (1991); Gao and Lemos (2008) and on light propagation in such joined spacetimes Wang et al. (2020). In this and the next section, we mainly give a relatively complete but still brief review of these results in preparation for the later image analysis.
@@
By the Israel junction conditions Israel (1966), the induced metric on the matching hypersurface $\Sigma$ must be continuous, $h_{ab}]=0$, where $X]=X_{+}-X_{-}$, whereas its normal derivative may be discontinuous. The resulting jump in the extrinsic curvature, $K_{ab}]$, encodes a surface stress-energy tensor localized on $\Sigma$.
@@
So far we have only required continuity of the induced metric along the shell. The expressions in Eq. (7) are computed separately in the two bulks and, a priori, need not represent the same geometric vector on $\Sigma$. However, if the unit normals from the two sides are not identified as a single normal field on $\Sigma$, the jump $K_{ab}]$ and the associated distributional definitions of the connection and curvature become ambiguous. This issue is more pronounced when three or more bulk spacetimes are glued along a single hypersurface. See Ref. Shen et al. (2024) and Appendix A for a detailed discussion. To ensure a well-defined differentiable structure and distributional curvature at the shell, we therefore identify the unit normals on the two sides at $\Sigma$:
@@
We consider a pressureless thin shell with surface stress-energy tensor $S^{i}{}_{j}=\mathrm{diag}(-\sigma,0,0)$. From the Israel conditions Israel (1966)
@@
Eq. (16) shows that $m_{\mathrm{sh}}$ is constant. For a dust shell with surface density $\sigma$, the conserved rest mass $m_{\rm sh}$ sets the amount of matter localized on the junction. The mass difference $\Delta m\equiv m_{+}-m_{-}$ represents the total energy carried by the shell as seen from infinity, including kinetic and gravitational binding contributions. Using Eq. (15) with $m_{\mathrm{sh}}=\mathrm{const}$ yields Wang et al. (2020)
@@
Figure 1: Schematic effective potential $V(R)$ for the thin spherical shell. The solid black curve shows $V(R)$; the horizontal dashed lines in red, blue, and orange mark $e^{2}$ corresponding to the cases $e^{2}\!<\!1$, $e^{2}\!=\!1$, and $e^{2}\!>\!1$, respectively. For $e^{2}<1$, the shell undergoes bound motion: it can reach at most the turning point $R_{\max}$, i.e., the intersection of the red dashed line with the black curve $V(R)$.
@@
Requiring $\sigma\geq 0$ implies $e\geq 0$. For $R>0$ the effective potential is monotonically increasing and concave, i.e. $V^{\prime}(R)>0$ and $V^{\prime\prime}(R)<0$. Fig. 1 schematically illustrates $V(R)$. When $e>1$ the motion is unbound and the shell can expand to spatial infinity; when $e=1$ the motion is critical and the shell reaches infinity with vanishing asymptotic velocity; when $e<1$ there is a turning point $R_{\max}$ defined by $V(R_{\max})=e^{2}$. Consequently, a pressureless shell admits no static, stable configuration, and static stability therefore requires nonzero pressure. Furthermore, any static shell must satisfy the necessary geometric constraint $R\geq 3m_{-}$, i.e., it must lie outside the interior photon sphere Brady et al. (1991).
@@ III Null geodesics in the joined spacetime
$\sin\alpha\equiv\frac{rk^{\phi}}{\sqrt{\left(k^{r}\right)^{2}+\left(rk^{\phi}\right)^{2}}},$ (26)
@@
$\sin\alpha_{+}=\left[1+\frac{\left(s\bigl(\beta_{+}(R)\beta_{-}(R)-\dot{R}^{2}\bigr)\sqrt{\,1-\sin^{2}\alpha_{-}\,}-\dot{R}\,(\beta_{+}(R)-\beta_{-}(R))\sqrt{\,1-(1-f_{-}(R))\sin^{2}\alpha_{-}\,}\right)^{2}}{f_{-}(R)^{2}\sin^{2}\alpha_{-}}\right]^{-1/2}$ (27)
@@
$E_{+}=\sqrt{\frac{f_{+}(R)}{f_{-}(R)}}\,E_{-},\qquad b_{+}=\sqrt{\frac{f_{-}(R)}{f_{+}(R)}}\,b_{-},\qquad\frac{\sin\alpha_{+}}{\sin\alpha_{-}}=\sqrt{\frac{1+\frac{f_{-}(R)}{f_{+}(R)}\left(\frac{R^{2}}{b_{+}^{2}}-f_{+}(R)\right)}{1+\left(\frac{R^{2}}{b_{+}^{2}}-f_{+}(R)\right)}}\,.$ (28)
From the null constraint, we have $\frac{R^{2}}{b_{+}^{2}}-f_{+}(R)\geq 0$. Thus, for a static shell, a photon exiting the shell is redshifted and its “refraction angle” increases.
We now clarify what we mean by the “refraction angle”. We do not refer to the incident angle measured in an orthonormal frame comoving with the shell. Since the photon four momentum is continuous as a geometric vector, the direction measured in that local frame does not jump across the shell. Instead, our “refraction angle” is defined with respect to the static observers in the inner and outer Schwarzschild regions. Owing to the difference in the Killing fields across the shell, the static observers on either side are defined differently. As Eq. (11) shows, their orthonormal bases do not match at the shell. Consequently, observer-dependent quantities such as the incident and outgoing angles, as well as the frequency, undergo a change across the shell. From Eq. (28), we see that the incident and outgoing angles are generally different across the shell. To streamline notation and make contact with the optics analogy, we introduce $n_{\pm}$ through the Snell-like relation
$\frac{n_{-}}{n_{+}}=\frac{\sin\alpha_{+}}{\sin\alpha_{-}}.$ (29)
Here we can view $n_{\pm}$ as effective refractive indices on the inner and outer sides of the shell. Requiring a nonnegative surface energy density implies $\sin\alpha_{+}\geq\sin\alpha_{-}$, and hence $n_{-}\geq n_{+}$. The effect of refraction on the image will be clarified in the section dedicated to imaging. We note, however, that refractive index interpretations in gravitational settings have been discussed in the literature Ramezani-Aval (2024); Masghatian et al. (2025); Nouri-Zonoz et al. (2022), but there is still no unique, universally accepted definition. In our work, we do not attempt to define a refractive index in a fundamental sense. Instead, we introduce this notion simply as a convenient way to describe how the angle between the photon propagation direction and the radial direction changes when the light ray crosses the shell. Since our specific setting and the aspects we emphasize differ from those in the above studies, the resulting relations and their explicit forms may accordingly be different.
The changes in photon energy when light crosses the shell are encoded in the redshift factor, which directly influences the final image. For collapsing shells, see Ref. Koga et al. (2025) for a systematic treatment; here we briefly summarize and adapt the relevant results. Define $g=\omega_{\text{obs}}/\omega_{\text{em}}$. In a static, spherically symmetric background, frequencies are defined with respect to static observers, then
@@ IV Static Shell Imaging
Having completed the junction of two Schwarzschild spacetimes and the analysis of null geodesics, we now turn to the imaging features of the resulting joined geometry. We begin with a static spherical shell for two reasons. First, the static case is simpler and allows us to identify the concrete impact of the shell on the image and the underlying mechanism, thereby laying the groundwork for understanding the various effects present for a dynamical shell. Second, the geometry obtained by matching two spacetimes along a static thin shell is of practical interest in its own right. For example, as mentioned in the Introduction, static shells are sometimes employed as idealized models of dark matter distributions around black holes Adhikari et al. (2014); Qian et al. (2022).
@@
where $r_{h_{-}}$ is the event horizon radius of the inner Schwarzschild region. We adopt this model precisely because our interest lies in image features controlled by the photon sphere. For a geometrically thin and optically thin disk, the photon ring of the final image is largely insensitive to the detailed radial emissivity, and extending the disk to the inner Schwarzschild horizon ensures that all trajectories relevant to the ring are included. For further discussion and comparisons of accretion disk models, see our previous work Cao et al. (2024b, a); Li et al. (2025). Accounting for the redshift factor, the observed specific intensity reads
@@
The V-shaped structure in Fig. 3(b) directly reflects the shell’s refraction effect. As derived in Sec. III, rays refract when traversing the shell, altering their propagation direction. The left panel of Fig. 4 shows rays that intersect the disk twice. The shell refracts rays with small $b$ outward, which increases the radius $r$ of the second intersection with the disk. As a result, $r(b)$ becomes nonmonotonic near $b\approx 3\sqrt{3}$ and takes on a V-shape in the plot. Similar V-shaped features of the transfer function $r(b)$ have, to our knowledge, been reported mainly in spacetimes with naked singularities or wormholes Luo et al. (2024); Rueda and Contreras (2025); Tan and Lan (2025). In contrast, in standard black hole spacetimes without a refractive shell, $r(b)$ is typically monotonic. In our setup, we find that for suitable shell parameters, adding a refractive shell around a black hole can produce such a V-shaped structure in the transfer function. The same behavior occurs for rays that intersect the disk three times. This is the reason for the $\Lambda$-shaped peak in Fig. 3(a).
Figure 5: Observed intensity profiles $I_{\rm obs}(b)$ and transfer functions $r(b)$ with $m_{-}=0.1$, shown for (a) $R_{\rm sh}=5$, (b) $R_{\rm sh}=3.1$, (c) $R_{\rm sh}=2.9$, and (d) $R_{\rm sh}=2.1$. In the right panels, black, orange, and red curves correspond to rays intersecting the disk once, twice, and three times, respectively. Insets in the $r(b)$ panels provide zoomed-in views of the local features.
Figure 6: Black hole images corresponding to the profiles in Fig. 5, for a static thin shell with fixed interior mass $m_{-}=0.1$. Panels (a)–(d) show the results for spherical shells at radii $R_{\rm sh}=5$, $3.1$, $2.9$, and $2.1$, respectively. In panel (d), we use a different intensity scale in order to highlight the double photon ring structure.
@@
Figure 7: Observed intensity profiles $I_{\rm obs}(b)$ and transfer functions $r(b)$ with $m_{-}=0.8$, shown for (a) $R_{\rm sh}=5$, (b) $R_{\rm sh}=3.1$, (c) $R_{\rm sh}=2.9$, and (d) $R_{\rm sh}=2.6$. In the right panels, black, orange, and red curves correspond to rays intersecting the disk once, twice, and three times, respectively. Insets in the $r(b)$ panels provide zoomed-in views of the local features.
Figure 8: Black hole images corresponding to the profiles in Fig. 7, for a static thin shell with fixed interior mass $m_{-}=0.8$. Panels (a)–(d) show the results for spherical shells at radii $R_{\rm sh}=5$, $3.1$, $2.9$, and $2.6$, respectively.